OpenStudy (anonymous):

Which is the equation of a line parallel to the line with equation y = 3x - 4?

OpenStudy (curry):

y=3x -4 unless you have options

OpenStudy (curry):

and by the grade

OpenStudy (curry):

dont forgot medal

OpenStudy (anonymous):

the choices are x - 3y = 1 3x + y = 7 x + 3y = 4 3x - y = -2

OpenStudy (curry):

D

OpenStudy (curry):

dont frogot medal and waht grade are u in

OpenStudy (lalaly):

parallel lines hav the same slope, which is the coeffecient of x...just rearrange the equations makin y subject.. the answer is the las one >> y=3x+2

OpenStudy (anonymous):

y=3x+k where k is a parameter